sleepyfish:

It takes one week for a crew of workers to pave ⅗ kilometer of a road. At that rate, how long will it take to pave 1 kilometer? Write a multiplication equation and a division equation that represent the question and then answer the question. Show your reasoning.

Angle:

We can set up a proportion to solve this problem Have one side be our starting ratio: 7 days over 3/5 kilometers have our other side be our missing ratio: x days over 1 kilometer It will look something like this: \(\huge \frac{7 ~days}{3/5 ~kilometers} = \frac{x~days}{1 ~kilometer}\) or \(\huge\frac{7}{3/5} = \frac{x}{1}\) do you know how to solve a proportion that looks like this?